# Learning Discrete Concepts in Latent Hierarchical Models

### v1

@@ Discussion.
Parallel to Theorem A4 [39] for linear models, Theorem 3.5 acts as an oracle to reveal the minimal t-separation set’s cardinality between any two variable sets in discrete models beyond linearity. This enables us to infer the latent graph structure from only observed variables’ statistical information. To the best of our knowledge, Theorem 3.5 is the first to establish this connection and can be of independent interests for learning latent discrete models in future work. Although the computation of non-negative ranks can be expensive [40], existing work [41, 42] demonstrates that regular rank tests are decent substitutes, we observe in our synthetic data experiments (Section 4). Since our focus is on fundamental identifiability, we experiment with relatively small-scale graphs (Section 4) and leave more efficient approximation as future work.
We present the identification conditions for discrete models as follows (Condition 3.7).
@@
Condition 3.7 (Discrete Hierarchical Model Conditions).
@@
Condition 3.7-i is known as the faithfulness condition widely adopted for causal discovery [43, 24, 44, 18], which attributes statistical independence to graph structures rather than unlikely coincidence [45, 43]. In linear models, Dong et al. [20] introduce atomic covers (Definition 3.6) to represent a group of indistinguishable variables, due to their identical neighbors. In the discrete case, an atomic cover consists of indistinguishable latent states, which we merge into a single latent discrete variable (Condition 3.1-ii). Intuitively, we treat each state as a separate variable and merge those belonging to the same atomic cover at the end of the identification procedure. This handles discrete variables of arbitrary state numbers, in contrast with the binary or identical support assumptions in prior work [22, 23], which we use as an alternative condition in Theorem 3.9. Condition 3.7-ii requires each atomic cover to leave a sufficient footprint (i.e., sufficient pure children and neighbors). In contrast, existing work [23] assumes at least three pure children for each latent variable, amounting to six times more states. Condition 3.7-iii allows us to discover latent colliders with sufficiently large supports, admitting graphs more general than tree structures [21, 26, 22] (i.e., no colliders). Overall, our model encompasses a rich class of latent hierarchical structures more complex than tree structures and multi-level DAGs [23], as illustrated in Figure 2.
Following Dong et al. [20], we introduce the minimal-graph operator $\mathcal{O}_{\min}$ (Definition A9), which merges certain redundancy structures that rank information cannot distinguish.
@@
Suppose the causal model $\mathcal{G}$ satisfies Condition 3.1 and Condition 3.7 We can identify $\mathcal{G}$ up to the Markov equivalence class of $\mathcal{O}_{\min}(\mathcal{G})$.
@@
As discussed above, Theorem 3.5 gives a graph structure oracle equivalent to Theorem A4, which we leverage to prove Theorem 3.8. Besides the rank test, the major distinction between Theorem A4 and Theorem 3.5 is that the former returns the number of variables in the minimal t-separation set whereas the latter returns the number of states. Applying the search algorithm from Dong et al. [20] alongside our rank test from Theorem 3.5 to a discrete model $\mathcal{G}$ results in a graph $\tilde{\mathcal{G}}$. In $\tilde{\mathcal{G}}$, each latent variable $z$ in $\mathcal{G}$ is split into a set of variables $\tilde{z}^{(1)},\dots,\tilde{z}^{(|\text{Supp}(z)|)}$ as an atomic cover, with the set size equal to the state number of $z$. We can then reconstruct the original graph $\mathcal{G}$ from $\tilde{\mathcal{G}}$ by merging these atomic covers into discrete variables whose state counts reflect the cover sizes. We present the specific search algorithm in Algorithm 1.
With the same technique, we show in Theorem 3.9 that the identical support condition (e.g., binary latent variables) adopted in prior work [23, 22] can also give rise to identification under slightly different conditions. We introduce the skeleton operator $\mathcal{O}_{s}$ [19, 20] (Definition A12) that include edges between adjacent covers indistinguishable to rank information.
Theorem 3.9 (Discrete Hierarchical Identification on Identical Supports).
Suppose the causal model $\mathcal{G}$ satisfies Condition 3.1-i, Condition A15, and $\left\lvert\text{Supp}(z)\right\rvert=K\geq 2$ for all $z\in\mathbf{Z}$. We can identify $\mathcal{G}$ up to the Markov equivalence class of $\mathcal{O}_{s}(\mathcal{O}_{\min}(\mathcal{G}))$.
@@ 4 Synthetic Data Experiments
Results and discussion. We choose Kivva et al. [24] as our baseline because it is the only method we know designed to learn a non-parametric, discrete latent model from continuous observations. We evaluate both methods on graphs in Figure A2. Our method consistently achieves near-perfect scores, while the baseline, despite correctly identifying $\Gamma$ and directing edges among $\mathbf{d}$ components, cannot handle higher-level latent variables.
To verify Theorem 3.5, we evaluate Algorithm 1 and a baseline [20] on graphs satisfying the conditions on $\mathcal{G}$ (i.e., purely discrete models in Figure A3). Our method performs well on graphs that meet conditions of Theorem 3.9 and achieves decent scores on graphs that do not (Figure A3 (c) and (e)). The significant margins over the baseline validate Theorem 3.5 and Theorem 3.9.
@@ 5 Interpretations of Latent Diffusion
Figure 4: Hierarchical Concept Ordering. We inject concepts of distinct abstraction levels into the generating process at different time steps. In the top row, the concept injection follows the hierarchical order, which renders injected concepts faithfully. The bottom row reverses the hierarchical order and cannot incorporate concepts properly. More examples in Figure A5.
@@
We connect LD training and estimating latent variables in the hierarchical model in an intuitive sense. Our theory focuses on the fundamental conditions on the data-generating process and do not directly translate to the guarantees for LD, which would require nontrivial assumptions on LD and we leave as future work. That said, our conditions naturally give implications on the algorithm design. For instance, a sparsity constraint on the decoding model may facilitate the identification condition that variables influence each other sparsely (e.g., pure children in Condition 3.7). In Section 6.2, we show such sparsity constraints are beneficial for concept extraction. We hope that our new perspective can provide more novel insights into advancing practical algorithms.
@@ 6 Real-world Experiments
6.1 Visualizing Hierarchical Concept Structures
Figure 5: Semantic latent space in the diffusion model. We modify the diffusion model’s representation (UNet encoder’s output) along principal directions at diffusion time steps $T$ and $0.6T$. Structure changes indicates the semantic property of the representation and manipulation at the early time $T$ induces global concept shifts. More examples in Figure A6.
In this section, we show that latent representations at different diffusion steps correspond to different levels of the hierarchical causal model. We select concept pairs, each with higher-level and lower-level concepts. For example, in (“sketch,” “wide eyes”), “sketch” is more global, while “wide eyes” is more local. We alter the text prompt during diffusion generation for concept injection, appending “in a sketch style” to inject “sketch” (see Appendix A5 for prompts). In Figure 4, global concepts are successfully injected at early diffusion steps and local ones at late steps (top row). Reversing this order fails, as shown in the bottom row. For example, injecting “sketch” early and “wide eyes” late renders both correctly, but the global concept “sketch” is absent under the reverse injection order. This supports our theory that concepts are hierarchically organized, with higher-level concepts related to earlier diffusion steps.
Further, we support our interpretations in Section 5 that diffusion representation can be viewed as concept embeddings, and it corresponds to high-level concepts for high noise levels. Following Park et al. [52], we modify the diffusion representation along certain directions found unsupervisedly. We can observe that this manipulation gives rise to semantic concept changes rather than entangled corruption Figure 5. Editing the latent representation at early steps corresponds to shifting global concepts. In Figure 5, the latent representation in earlier steps (step $T$) determines breeds (the top row), species (the middle row), and gender (the bottom row). In contrast, the latent representation in later steps (step $0.6T$) correlates with the dog collar, cat eyes, and shirt patterns. Implementation details and additional results are provided in Appendix A5.
@@ 6.2 Causal Sparsity for Concept Extraction
Figure 6: Concepts have varying levels of sparsity. We show that concepts of various abstraction levels correspond to different sparsity levels. For instance, bright weather is appropriately conveyed by a rank-$2$ LoRA and higher-rank LoRAs alter the background. Inadequate ranks fail to capture the concept faithfully and unnecessary ranks entangle the target concept with other attributes.
(a) CLIP scores.
(b) LPIPS scores.
Figure 7: CLIP/LPIPS evaluation. We evaluate our approach and baselines at individual rank constraints. A high CLIP score is favorable as it indicates semantic alignment. A low LPIPS score is more favorable as it indicates minimal excessive changes. We compare our method “sparse” with the optimal fixed rank setting on each concept. For instance, “castle opt: rank4_scale3” indicates that the optimal setting for the concept “castle” is the LoRA of rank $4$ and scale $3$. With a adaptive rank selection, our approach outperforms or keeps up with the optimal fixed setting across different concepts. We repeat each training over three random seeds.
Recent work [55] shows that concepts can be extracted as low-rank parameter subspaces of LD models via LoRA [56]. This low-rankness limits the complexity of text-induced changes, resembling sparse influences from latent concepts to their descendants. Our theory suggests that different levels of concepts may require varying sparsity levels to capture. Figure 6 shows that indeed concepts at different abstraction levels have desirable representations at different ranks. For instance, the concept of bright weather is appropriately conveyed by a rank-$2$ LoRA and higher-rank LoRAs alter the background. The same observation occurs to other concepts, where inadequate ranks fail to capture the concept faithfully and unnecessary ranks entangle the target concept with other attributes.
Motivated by this, we design an adaptive sparsity selection mechanism for capturing concepts at different levels. Specifically, inspired by Ding et al. [57], we implement a sparsity constraint on the LoRA dimensionality for the model to select the LoRA rank at each module automatically (see Appendix A5 for details). Figure 7 presents the CLIP and LPIPS evaluation for the baseline and our approach, where the CLIP score evaluates the alignment between the image and the target description and the LPIPS score measures the structure change between the edited image and the original image. We can observe that under the sparsity constraint, our approach attains the highest CLIP score and the lowest LPIPS score when compared with the baselines of several ranks, indicating a higher level of alignment and a lower level of undesirable entanglement.
@@ 7 Conclusion
In this work, we cast the task of learning concepts as the identification problem of a discrete latent hierarchical model. Our theory provides conditions to guarantee the recoverability of discrete concepts. Further, we discuss and evaluate our framework’s implications on the working mechanism of diffusion models. Limitations: Although our theoretical framework provides a lens to interpret diffusion models, our conditions do not directly guarantee diffusion’s success, which would require nontrivial assumptions on the diffusion model. Also, Algorithm 1 can be expensive for large graphs, as the probability table’s size grows linearly with the number of discrete states. We leaving giving guarantees to diffusion models and more efficient graph learning algorithm as future work.
@@
In recent years, a significant strand of research has focused on employing labeled data to learn concepts in generative models’ latent space for image editing and manipulation [1, 2, 3, 4, 5, 6]. Concurrently, another independent research trajectory has been exploring unsupervised concept discovery and its potential to learn more compositional and transferable models, as shown in Burgess et al. [7], Locatello et al. [8], Du et al. [9, 10], Liu et al. [11]. These prior works focus on the empirical methodological development of concept learning by proposing novel neural network architectures and training objectives, with limited discussion on the theoretical aspect. In contrast, our work investigates the theoretical foundation of concept learning. Specifically, we formulate concept learning as an identification problem for a discrete latent hierarchical model and provide conditions under which extracting concepts is possible. Thus, the existing work and our work can be viewed as two complementary lines of research for concept learning.
@@
Complex real-world data distributions often possess a hierarchical structure among their underlying latent variables. On the theoretical front, Xie et al. [18], Huang et al. [19], Dong et al. [20] investigate identification conditions of latent hierarchical structures under the assumption that the latent variables are continuous and influence each other through linear functions. Kong et al. [58] extends the functional class to the nonlinear case over continuous variables. Pearl [21], Zhang [26], Choi et al. [22], Gu and Dunson [23] study fully discrete cases and thus fall short of modeling the continuous observed variables like images. Specifically, Pearl [21], Zhang [26], Choi et al. [22] focus on the latent trees in which every pair of variables is connected through exactly one undirected path. Gu and Dunson [23] assume a multi-level DAG [59] in which variables can be partitioned into disjoint groups (i.e., levels), such that all edges are between adjacent levels, with the observed variables as the bottom level (i.e., leaf nodes). In contrast, we show that we can not only extract discrete components from continuous observed variables but also uncover higher-level concepts and their interactions. Our graphical conditions admit multiple paths within each pair of latent variables, flexible hierarchical structures that are not necessarily multi-level, and flat structures in which all latent variables are adjacent to observed variables [24]. On the empirical side, prior work [60] improves the inference model of vanilla VAEs by combining bottom-up data-dependent likelihood terms with prior generative distribution parameters. Zhao et al. [61] assign more expressive (deeper) neural modules to higher-level variables to learn a more disentangled generative model. Li et al. [62] present a VAE architecture and clustering approaches to empirical estimate latent tree structures. Leeb et al. [63] propose to feed latent variable partitions into different decoder neural network layers and remove the prior regularization term to enable high-quality generation. Like our work, Ross and Doshi-Velez [64] consider discrete latent variables. However, their focus is on empirical evaluation benchmarks and metrics, without touching on the theoretical formulation of this task. Unlike these efforts, our work concentrates on the formalization of the data-generating process and the theoretical understanding. Thus, these two lines complement each other.
@@
The estimate $\hat{\mathbf{d}}$ and the true variable $\mathbf{d}$ are related through the map $[\hat{\mathbf{d}},\hat{\mathbf{c}}]=\hat{g}^{-1}\circ g(\mathbf{d},\mathbf{c})$. In the following, we show that the induced relation between $\mathbf{d}$ and $\hat{\mathbf{d}}$ is invertible under Condition 3.3-ii. The estimated generating process respects the conditions on the true generating process.
@@
Suppose that $k$ corresponds to two distinct states $\hat{k}_{1}$ and $\hat{k}_{2}$. That is, there exist $\mathbf{c}_{1},\mathbf{c}_{2}\in\mathcal{C}$ and $\hat{\mathbf{c}}_{1},\hat{\mathbf{c}}_{2}\in\hat{\mathcal{C}}$, such that $\hat{g}^{-1}\circ g(k,\mathbf{c}_{1})=[\hat{k}_{1},\hat{\mathbf{c}}_{1}]$ and $\hat{g}^{-1}\circ g(k,\mathbf{c}_{2})=[\hat{k}_{2},\hat{\mathbf{c}}_{2}]$. On one hand, As $g(k,\cdot)$ is a continuous function and $\mathcal{C}$ is connected, the image $\mathcal{I}(k):=g(k,\mathcal{C})$ is a connected set. On the other hand, ^$\hat{\mathcal{I}}(k_{1}):=\hat{g}(\hat{k}_{1},\hat{\mathcal{C}})$ and ^$\hat{\mathcal{I}}(\hat{k}_{2}):=\hat{g}(\hat{k}_{2},\hat{\mathcal{C}})$ are two separate sets due to the invertibility and continuity of $\hat{g}$ and the openness of $\hat{\mathcal{C}}$. To see this, invertibility implies that ^$\hat{\mathcal{I}}(k_{1})$ and ^$\hat{\mathcal{I}}(k_{2})$ are disjoint. The fact that $\hat{g}$ is continuous over $\hat{\mathbf{c}}$ and has a continuous inverse over $\hat{\mathbf{c}}$ implies that ^$\hat{\mathcal{I}}(k_{1})$ and ^$\hat{\mathcal{I}}(k_{2})$ preserve the openness $\mathcal{C}$. The space formed by two disjoint open subspaces is disconnected. Since $k$ corresponds to $\hat{k}_{1}$ and $\hat{k}_{2}$, it follows that $\mathcal{I}(k)=\hat{\mathcal{I}}_{1}\cup\hat{\mathcal{I}}_{2}$ where ^$\hat{\mathcal{I}}_{1}$ and ^$\hat{\mathcal{I}}_{2}$ are nonempty subsets of ^$\hat{\mathcal{I}}_{\hat{k}_{1}}$ and ^$\hat{\mathcal{I}}_{\hat{k}_{2}}$ respectively and inherit their separability. As $\mathcal{I}(k)$ is a union of two nonempty separated sets, it is disconnected. This contradicts the connectedness of $\mathcal{I}(k)$. Therefore, for each state $k\in\Omega^{(d)}$, $k$ corresponds to at most one state $\hat{k}\in\hat{\Omega}^{(d)}$.
@@
Under Condition 3.1 and Condition 3.3-iii, on can reconstruct the bipartite graph $\Gamma$ between $\mathbf{d}$ and $\mathbf{x}$, and the joint distribution ${\mathbb{P}}\left(d_{1}=k_{1},\dots,d_{n_{d}}=k_{n_{d}}\right)$ from ${\mathbb{P}}\left(\mathbf{x}\right)$ and $\text{MixOracle}(\mathbf{x})$.

### v2

@@ Discussion.
Parallel to Theorem A2 [49] for linear models, Theorem 4.8 acts as an oracle to reveal the minimal t-separation set’s cardinality between any two variable sets in discrete models beyond linearity. This enables us to infer the latent graph structure from only the observed variables’ statistical information. To the best of our knowledge, Theorem 4.8 is the first to establish this connection and can be of independent interest for learning latent discrete models in future work. Although the computation of non-negative ranks can be expensive [50], existing work [52, 53] demonstrates that regular rank tests are decent substitutes, we observe in our synthetic data experiments (Section 5).
We present the identification conditions for discrete models as follows (Condition 4.10).
@@
Condition 4.10 (Discrete Hierarchical Model Conditions).
@@
Condition 4.10-i is known as the faithfulness condition widely adopted for causal discovery [51, 24, 54, 18], which attributes statistical independence to graph structures rather than unlikely coincidence [55, 51]. In linear models, Dong et al. [20] introduce atomic covers (Definition 4.9) to represent a group of indistinguishable variables. In the discrete case, an atomic cover consists of indistinguishable latent states, which we merge into a single latent discrete variable (Condition 4.1-ii). Intuitively, we treat each state as a separate variable and merge those belonging to the same atomic cover at the end of the identification procedure. This handles discrete variables of arbitrary state numbers, in contrast with the binary or identical support assumptions [22, 23], which we use as an alternative condition in Theorem A12. Condition 4.10-ii requires each atomic cover to possess sufficiently many children and neighbors to preserve its influence while avoiding problematic triangle structures to ensure the uniqueness of its influence. In contrast, existing work [23] assumes at least three pure children for each latent variable, amounting to six times more states. Condition 4.10-iii ensures adequate side information (large $|\mathbf{A}|$) to discover latent colliders $\mathbf{C}$, admitting graphs more general than tree structures [21, 26, 22] (i.e., no colliders). Overall, our model encompasses a rich class of latent structures more complex than tree structures and multi-level DAGs [23] (Figure 2).
Following Dong et al. [20], we introduce the minimal-graph operator $\mathcal{O}_{\min}$ (Definition 4.11 and Figure A1), which merges certain redundancy structures that rank information cannot distinguish.
Definition 4.11 (Minimal-graph Operator [19, 20]).
We can merge atomic covers $\mathbf{L}$ into $\mathbf{P}$ in $\mathcal{G}$ if (i) $\mathbf{L}$ is a pure child of $\mathbf{P}$, (ii) all elements of $L$ and $P$ are latent and $|\text{Supp}(\mathbf{L})|=|\text{Supp}(\mathbf{P})|$, and (iii) the pure children of $\mathbf{L}$ form a single atomic cover, or the siblings of $\mathbf{L}$ form a single atomic cover. We denote such an operator as the minimal-graph operator $\mathcal{O}_{\min}(\mathcal{G})$.
@@
Suppose the causal model $\mathcal{G}$ satisfies Condition 4.1 and Condition 4.10 We can identify $\mathcal{G}$ up to the Markov equivalence class of $\mathcal{O}_{\min}(\mathcal{G})$.
@@
As discussed above, Theorem 4.8 gives a graph structure oracle equivalent to Theorem A2, which we leverage to prove Theorem 4.12. Besides the rank test, the major distinction between Theorem A2 and Theorem 4.8 is that the former returns the number of variables in the minimal t-separation set whereas the latter returns the number of states. Applying the search algorithm from Dong et al. [20] alongside our rank test from Theorem 4.8 to a discrete model $\mathcal{G}$ results in a graph $\tilde{\mathcal{G}}$. In $\tilde{\mathcal{G}}$, each latent variable $z$ in $\mathcal{G}$ is split into a set of variables $\tilde{z}^{(1)},\dots,\tilde{z}^{(|\text{Supp}(z)|)}$ as an atomic cover, with the set size equal to the state number of $z$. We can then reconstruct the original graph $\mathcal{G}$ from $\tilde{\mathcal{G}}$ by merging these atomic covers into discrete variables. We present our algorithm in Algorithm 1 and highlight the differences from that in Dong et al. [20].
Our techniques can also utilize the identical support condition (e.g., binary latent variables) [23, 22] for identification under slightly different conditions. We present the results in Theorem A12.
@@ 5 Synthetic Data Experiments
Results and discussion. We choose Kivva et al. [24] as our baseline because it is the only method we know designed to learn a non-parametric, discrete latent model from continuous observations. We evaluate both methods on graphs in Figure A2. As shown in Table 1 and Table 2, our method consistently achieves near-perfect scores, while the baseline, despite correctly identifying $\Gamma$ and directing edges among $\mathbf{d}$ components, cannot handle higher-level latent variables.
To verify Theorem 4.8, we evaluate Algorithm 1 and a baseline [20] on graphs satisfying the conditions on $\mathcal{G}$ (i.e., purely discrete models in Figure A3). Our method performs well on graphs that meet conditions of Theorem A12 and achieves decent scores on graphs that do not (Figure A3 (c) and (e)). The significant margins over the baseline validate Theorem 4.8 and Theorem A12.
@@ 6 Interpretations of Latent Diffusion
We connect LD training and estimating latent variables in the hierarchical model in an intuitive sense. Our theory focuses on the fundamental conditions of the data-generating process and does not directly translate to guarantees for LD. That said, our conditions naturally have implications on the algorithm design. For instance, a sparsity constraint on the decoding model may facilitate the identification condition that variables influence each other sparsely (e.g., pure children in Condition 4.10). In Section 7.3, we show such constraints are beneficial for concept extraction. We hope that our new perspective can provide more novel insights into advancing practical algorithms.
@@ 7 Real-world Experiments
7.1 Discovering Hierarchical Concept Structures from Diffusion Models
Figure 4: Recovering concepts and their relationships from LD. (a) The final recovered concept graph among concepts “dog”, “tree”, “eyes”, “ears”, “branch”, and “leaf”. (b) Identifying causal links through “interventions”. For example, we compare two prompts that vary in “dog”: “a dog with wide eyes and a wilting tree with short branches, in a cartoon style” and “a big dog with wide eyes and a wilting tree with short branches, in a cartoon style”. We observe significant changes in “eyes” but not in “branch”, indicating a causal link between “dog” and “eyes” but not between “dog” and “branch”. (c) Identifying concept levels by the last effective diffusion step. For example, we use the base prompt “a tree with long branches, in a cartoon style” and prepend “dog” at steps 0, 5, and 15. Only injecting “dog” at step 0 works. Similarly, injecting “wide eyes” works at both steps 0 and 5, indicating that “dog” is a higher-level concept than “eyes”.
In Figure 4, we extract concepts and their relationships from LD through our hierarchical model interpretation. Our recovery involves two stages: determining the concept level and identifying causal links. We add a textual concept, like “dog”, into the prompt and identify the latest diffusion step that would render this concept properly. If “dog” appears in the image only when added at step 0 and “eye” appears when added from step 5, it indicates that “dog” is a higher-level concept than “eyes”. After determining the levels of concepts, we intervene on a high-level concept and observe changes in low-level ones. No significant changes indicate no direct causal relationship. We explore the relationships among the concepts “dog”, “tree”, “eyes”, “ears”, “branch”, and “leaf”. Figure 4 presents the final recovered graph and intermediate results. See Section A5.3 for more investigation.
7.2 Diffusion Representation as Concept Embeddings
We support our interpretations in Section 6 that diffusion representation can be viewed as concept embeddings, and it corresponds to high-level concepts for high noise levels. Following Park et al. [62], we modify the diffusion representation along certain directions found unsupervisedly. We can observe that this manipulation gives rise to semantic concept changes rather than entangled corruption Figure 5. Editing the latent representation at early steps corresponds to shifting global concepts.
Figure 5: Semantic latent space. We modify the diffusion model’s representation (UNet encoder’s output) along principal directions at steps $T$ and $0.6T$. Structure changes indicate the semantics of the representation and manipulation at the early time $T$ induces global shifts. See more examples in Figure A9.
In Figure 5, the latent representation in earlier steps (step $T$) determines breeds (the top row), species (the middle row), and gender (the bottom row). In contrast, the latent representation in later steps (step $0.6T$) correlates with the dog collar, cat eyes, and shirt patterns. Implementation details and additional results are provided in Appendix A5.
@@ 7.3 Causal Sparsity for Concept Extraction
Recent work [65] shows that concepts can be extracted as low-rank parameter subspaces of LD models via LoRA [66]. This low-rankness limits the complexity of text-induced changes, resembling sparse influences from latent concepts to their descendants. Our theory suggests that different levels of concepts may require varying sparsity levels to capture. We present empirical evidence in Section A5.4. Motivated by this, we design an adaptive sparsity selection mechanism for capturing concepts at different levels. Inspired by Ding et al. [67], we implement a sparsity constraint on the LoRA dimensionality for the model to select the LoRA rank at each module automatically, benefiting concept extraction (see Appendix A5.4).
@@ 8 Conclusion
In this work, we cast the task of learning concepts as the identification problem of a discrete latent hierarchical model. Our theory provides conditions to guarantee the recoverability of discrete concepts. Limitations: Although our theoretical framework provides a lens for interpretation, our conditions do not directly guarantee diffusion’s success, which would require nontrivial assumptions. Also, Algorithm 1 can be expensive for large graphs due to the dependency on the state count. We leave giving guarantees to diffusion models and efficient graph learning algorithms as future work.
Acknowledgments.
We thank the anonymous reviewers for their valuable insights and recommendations, which have greatly improved our work. The work of L. Kong and Y. Chi is supported in part by NSF DMS-2134080. This material is based upon work supported by NSF Award No. 2229881, AI Institute for Societal Decision Making (AI-SDM), the National Institutes of Health (NIH) under Contract R01HL159805, and grants from Salesforce, Apple Inc., Quris AI, and Florin Court Capital. This research has been graciously funded by the National Science Foundation (NSF) CNS2414087, NSF BCS2040381, NSF IIS2123952, NSF IIS1955532, NSF IIS2123952; NSF IIS2311990; the National Institutes of Health (NIH) R01GM140467; the National Geospatial Intelligence Agency (NGA) HM04762010002; the Semiconductor Research Corporation (SRC) AIHW award 2024AH3210; the National Institute of General Medical Sciences (NIGMS) R01GM140467; and the Defense Advanced Research Projects Agency (DARPA) ECOLE HR00112390063. Any opinions, findings, and conclusions or recommendations expressed in this publication are those of the author(s) and do not necessarily reflect the views of the National Science Foundation, the National Institutes of Health, the National Geospatial Intelligence Agency, the Semiconductor Research Corporation, the National Institute of General Medical Sciences, and the Defense Advanced Research Projects Agency.
@@ Acknowledgments.
In recent years, a significant strand of research has focused on employing labeled data to learn concepts in generative models’ latent space for image editing and manipulation [1, 2, 3, 4, 5, 6]. Concurrently, another independent research trajectory has been exploring unsupervised concept discovery and its potential to learn more compositional and transferable models, as shown in Burgess et al. [7], Locatello et al. [8], Du et al. [9, 10], Liu et al. [11]. These prior works focus on the empirical methodological development of concept learning by proposing novel neural network architectures and training objectives, with limited discussion on the theoretical aspect. In contrast, our work investigates the theoretical foundation of concept learning. Specifically, we formulate concept learning as an identification problem for a discrete latent hierarchical model and provide conditions under which extracting concepts is possible. Thus, the existing work and our work can be viewed as two complementary lines of research for concept learning. Concurrently, a plethora of work has been dedicated to extracting interpretable concepts from high-dimensional data such as images. Concept-bottleneck [12] first predicts a set of human-annotated concepts as an intermediate stage and then predicts the task labels from these intermediate concepts. This paradigm has attracted a large amount of follow-up work [13, 29, 30, 31, 32, 33]. A recent surge of pre-trained multimodal models (e.g., CLIP [17]) can explain the image concepts through text directly [14, 15, 16]. In contrast with these successes, our work focuses on the formulation of concept learning and theoretical guarantees.
@@
Complex real-world data distributions often possess a hierarchical structure among their underlying latent variables. On the theoretical front, Xie et al. [18], Huang et al. [19], Dong et al. [20] investigate identification conditions of latent hierarchical structures under the assumption that the latent variables are continuous and influence each other through linear functions. Kong et al. [34] extends the functional class to the nonlinear case over continuous variables. Pearl [21], Zhang [26], Choi et al. [22], Gu and Dunson [23] study fully discrete cases and thus fall short of modeling the continuous observed variables like images. Specifically, Pearl [21], Zhang [26], Choi et al. [22] focus on the latent trees in which every pair of variables is connected through exactly one undirected path. Gu and Dunson [23] assume a multi-level DAG [68] in which variables can be partitioned into disjoint groups (i.e., levels), such that all edges are between adjacent levels, with the observed variables as the bottom level (i.e., leaf nodes). In contrast, we show that we can not only extract discrete components from continuous observed variables but also uncover higher-level concepts and their interactions. Our graphical conditions admit multiple paths within each pair of latent variables, flexible hierarchical structures that are not necessarily multi-level, and flat structures in which all latent variables are adjacent to observed variables [24]. On the empirical side, prior work [69] improves the inference model of vanilla VAEs by combining bottom-up data-dependent likelihood terms with prior generative distribution parameters. Zhao et al. [70] assign more expressive (deeper) neural modules to higher-level variables to learn a more disentangled generative model. Li et al. [71] present a VAE/clustering approach to empirically estimating latent tree structures. Leeb et al. [72] propose to feed latent variable partitions into different decoder neural network layers and remove the prior regularization term to enable high-quality generation. Like our work, Ross and Doshi-Velez [73] consider discrete latent variables. However, their focus is on empirical evaluation benchmarks and metrics, without touching on the theoretical formulation of this task. Unlike these efforts, our work concentrates on the formalization of the data-generating process and the theoretical understanding. Thus, these two lines complement each other.
@@
The estimate $\hat{\mathbf{d}}$ and the true variable $\mathbf{d}$ are related through the map $[\hat{\mathbf{d}},\hat{\mathbf{c}}]=\hat{g}^{-1}\circ g(\mathbf{d},\mathbf{c})$. In the following, we show that the induced relation between $\mathbf{d}$ and $\hat{\mathbf{d}}$ is invertible under Condition 4.3-ii. The estimated generating process respects the conditions on the true generating process.
@@
Suppose that $k$ corresponds to two distinct states $\hat{k}_{1}$ and $\hat{k}_{2}$. That is, there exist $\mathbf{c}_{1},\mathbf{c}_{2}\in\mathcal{C}$ and $\hat{\mathbf{c}}_{1},\hat{\mathbf{c}}_{2}\in\hat{\mathcal{C}}$, such that $\hat{g}^{-1}\circ g(k,\mathbf{c}_{1})=[\hat{k}_{1},\hat{\mathbf{c}}_{1}]$ and $\hat{g}^{-1}\circ g(k,\mathbf{c}_{2})=[\hat{k}_{2},\hat{\mathbf{c}}_{2}]$. On one hand, As $g(k,\cdot)$ is a continuous function and $\mathcal{C}$ is connected, the image $\mathcal{I}(k):=g(k,\mathcal{C})$ is a connected set. On the other hand, ^$\hat{\mathcal{I}}(k_{1}):=\hat{g}(\hat{k}_{1},\hat{\mathcal{C}})$ and ^$\hat{\mathcal{I}}(\hat{k}_{2}):=\hat{g}(\hat{k}_{2},\hat{\mathcal{C}})$ are two separate sets due to the invertibility and continuity of $\hat{g}$ and the closed-ness of $\hat{\mathcal{C}}$. To see this, invertibility implies that ^$\hat{\mathcal{I}}(k_{1})$ and ^$\hat{\mathcal{I}}(k_{2})$ are disjoint. The fact that $\hat{g}$ is continuous over $\hat{\mathbf{c}}$ and has a continuous inverse over $\hat{\mathbf{c}}$ implies that ^$\hat{\mathcal{I}}(k_{1})$ and ^$\hat{\mathcal{I}}(k_{2})$ preserve the closed-ness of $\mathcal{C}$. The space formed by two disjoint closed subspaces is disconnected. Since $k$ corresponds to $\hat{k}_{1}$ and $\hat{k}_{2}$, it follows that $\mathcal{I}(k)=\hat{\mathcal{I}}_{1}\cup\hat{\mathcal{I}}_{2}$ where ^$\hat{\mathcal{I}}_{1}$ and ^$\hat{\mathcal{I}}_{2}$ are nonempty subsets of ^$\hat{\mathcal{I}}_{\hat{k}_{1}}$ and ^$\hat{\mathcal{I}}_{\hat{k}_{2}}$ respectively and inherit their separability. As $\mathcal{I}(k)$ is a union of two nonempty separated sets, it is disconnected. This contradicts the connectedness of $\mathcal{I}(k)$. Therefore, for each state $k\in\Omega^{(d)}$, $k$ corresponds to at most one state $\hat{k}\in\hat{\Omega}^{(d)}$.
@@
Under Condition 4.1 and Condition 4.3-iii, on can reconstruct the bipartite graph $\Gamma$ between $\mathbf{d}$ and $\mathbf{x}$, and the joint distribution ${\mathbb{P}}\left(d_{1}=k_{1},\dots,d_{n_{d}}=k_{n_{d}}\right)$ from ${\mathbb{P}}\left(\mathbf{x}\right)$ and $\text{MixOracle}(\mathbf{x})$.